# SpikeDet: Better Firing Patterns for Accurate and Energy-Efficient Object Detection with Spiking Neuron Networks

### v3

@@ V-A Experimental Setup
For I-LIF neurons, the membrane time constant $\tau$ is initialized to 0.25. The models are trained on 4 NVIDIA A6000 GPUs, using the SGD optimizer with a learning rate of $0.01$. During post-processing, we use NMS, and the top 100 boxes are selected for evaluation. Additionally, we use Complete IoU (CIOU) [52] and Distribution Focal Loss (DFL) [53] for bounding box regression, and Binary Cross Entropy (BCE) [2] for classification loss. For multiscale evaluation, following [54], we employ both original and horizontally flipped images at scale factors of 0.6, 1.0, 1.2, 1.5, and 1.8.
@@
The COCO 2017 is a large-scale object detection benchmark with 118,287 training and 5000 validation images, which we use for training and testing, respectively. Objects from 80 categories are annotated with their classes and locations. On this dataset, we train for 400 epochs with a batch size of 64 at a resolution of 640$\times$640, employing mosaic [55] and mixup [56] data augmentation for training and testing.
@@
Following [39, 55, 54, 4], we report Average Precision (AP) as the primary metric, along with AP at IoU thresholds of 0.5 and 0.75, as well as AP for small, medium, and large objects. The firing rate, which measures neuronal activity, is another critical metric for evaluating SNNs. It is calculated as the average ratio of neuron spikes to the total number of neurons across all time steps. Additionally, we report the power consumption of the models, and the calculation method is shown in III-B.
V-B Ablation Studies
@@
TABLE I: The ablation study on MDSNet. The specific configuration of the proposed MDSNet is detailed in Appendix B. and indicate MDSNet with MDS removed and replaced by membrane-based shortcuts, respectively.
Backbone Depth AP $\text{AP}_{50}$ Param (M) Firing Rate(%) Power (mJ)
EMS-ResNet [15] 34 43.2 59.5 22.9 14.6 28.7
SEW-ResNet [34] 34 43.2 59.7 22.9 12.2 151.9
MS-ResNet [25] 34 43.8 60.5 22.9 12.2 29.0
S-Backbone [20] 47 44.0 60.6 24.2 12.2 112.0
MDSNet 34 44.8 61.2 22.0 11.8 22.9
MDSNet 34 43.7 60.5 21.9 11.4 22.1
MDSNet 34 44.2 60.6 21.9 12.0 23.2
MDSNet 10 41.5 58.2 14.3 12.9 20.2
MDSNet 18 43.1 59.5 17.5 12.4 21.2
MDSNet 104 46.5 63.2 48.2 9.1 28.8
V-B1 The Effectiveness of MDSNet
To demonstrate the feature extraction capabilities of MDSNet, we compare it with the most advanced SNN backbones available. For fair comparison, we construct models using the backbone networks paired with identical neck and detection head architectures. The comparison results are shown in rows 1 to 5 of Table I. These results show that MDSNet34 significantly outperforms other models in AP and $\text{AP}_{50}$ while achieving a lower firing rate and power consumption. Additionally, we conduct comparative experiments by replacing MDS with membrane-based shortcuts and removing MDS. As shown in rows 5 to 7 of Table I, replacing MDS with membrane-based shortcuts demonstrates the effectiveness of MDS, while removing MDS validates the importance of residual learning structures. All these modifications result in performance degradation.
V-B2 Impact of Model Depth
In Section IV-C, we theoretically demonstrate that the MDSNet could overcome the issues of gradient vanishing or exploding, enabling the realization of deep networks. In this section, we further validate this through experiments. As shown in rows 5 and 8 to 10 of Table I, the feature extraction capability of MDSNet improves as the network depth increases. Notably, this is the first SNN backbone with over 100 layers trained on object detection datasets.
@@ V-B2 Impact of Model Depth
Fusion Method Fusion Directions AP $\text{AP}_{50}$ Param (M) Firing Rate(%) Power (mJ)
None 0 35.1 50.4 12.6 14.1 11.2
FPN [57] 1 38.9 56.4 14.9 14.2 15.3
PAN [58] 2 39.8 57.7 17.2 12.9 15.2
BiFPN [59] 4 38.7 56.2 12.0 14.6 12.9
SMFM 4 44.8 61.2 22.0 11.8 22.9
SMFM 1 41.3 58.6 21.1 13.1 25.3
SMFM 2 43.0 59.9 21.6 12.2 21.7
SMFM 6 44.2 60.7 22.1 11.3 22.1
V-B3 The Effectiveness of SMFM
We first compare the model without the feature fusion module with the one using SMFM. As shown in rows 1 and 5 of Table II, SMFM significantly improves AP by 9.7% with only slightly increased power consumption while also reducing firing rate. Furthermore, we compare our SMFM with the classical feature fusion modules of ANN detectors [57, 58, 59]. These methods are adapted for SNNs by replacing activations with I-LIF neurons. As shown in rows 2-5 of Table II, SMFM not only achieves superior performance but also effectively reduces firing rate with comparable power consumption. This demonstrates that our method effectively considers SNN characteristics and highlights the necessity of designing SNN-friendly feature fusion modules.
As shown in rows 5-8 of Table II, we compare the impact of different numbers of fusion directions on performance. To eliminate the effect of parameters, we adjust the width to maintain comparable parameters across models. Performance improves with more fusion directions when below 4, validating our analysis in IV-D. However, performance degrades beyond 4 directions, indicating that excessive complexity hinders optimization.
TABLE III: Comparison of different Fusion Blocks. The SF-Block1 refers to retaining only the first sub-block of SF-Block.
Fusion Block AP $\text{AP}_{50}$ Param (M) Firing Rate(%) Power (mJ)
7$\times$7-Conv 41.7 58.5 22.6 11.6 14.5
MS-Block 43.6 60.7 18.7 13.3 20.0
MDS-Block1 44.2 60.9 21.0 12.1 20.3
SF-Block1 43.8 60.9 20.3 12.0 20.9
SF-Block 44.8 61.2 22.0 11.8 22.9
V-B4 The Effectiveness of SF-Block
We replace the SF-Block in SMFM with other classical SNN-based blocks [25] and the simplest single convolution to demonstrate its effectiveness. As shown in Table III, SF-Block significantly outperforms other methods in both AP and firing rate, with only a slight increase in power consumption.
@@ V-B4 The Effectiveness of SF-Block
Model T×$\times$×D AP $\text{AP}_{50}$ Param (M) Firing Rate(%) Power (mJ)
1$\times$2 39.9 56.3 23.1 18.8 25.7
SpikeYOLO 1$\times$4 43.9 60.4 23.1 12.2 33.0
1$\times$6 45.2 61.9 23.1 9.0 36.4
1$\times$2 41.2 56.8 22.0 18.1 18.1
SpikeDet 1$\times$4 44.8 61.2 22.0 11.8 22.9
1$\times$6 45.6 62.3 22.0 8.8 25.4
V-B5 Impact of Time Step
Since the SpikeDet provides more stable firing patterns and overcomes local firing saturation problems, it should theoretically achieve comparable performance with fewer time steps. To demonstrate this, we conduct corresponding experiments. Since adjusting T has been demonstrated to be inefficient in [20], we control the overall time steps of neurons by adjusting D instead. As shown in Table IV, compared to SpikeYOLO [20], our method achieves good performance even in low time steps. Moreover, performance gains remain modest as the time steps increase, demonstrating stronger robustness to variations in time steps.
TABLE V: Performance Comparison with State-of-the-Art Models on COCO 2017 Validation Dataset. $\dagger$ indicates results from multiscale testing. The EMS-YOLO [15] is our reproduced version using I-LIF neurons for fair comparison.
Method Model AP AP50 AP75 APs APm APl Param (M) T×$\times$×D Firing Rate(%) Power (mJ)
ANNs SSD513 [60] 31.2 50.4 33.3 10.2 34.5 49.8 36.0 1 - 227.3
ATSS [23] 39.4 57.6 42.8 23.6 42.9 50.3 32.3 1 - 377.2
YOLOH [9] 40.3 60.9 43.5 22.9 44.4 54.0 46.0 1 - 195.5
DETR [41] 44.9 64.7 47.7 23.7 49.5 62.3 60.6 1 - 388.3
@@ V-B5 Impact of Time Step
Gold-YOLO-N [39] 39.6 55.7 - 19.7 44.1 57.0 5.6 1 - 27.6
Gold-YOLO-S [39] 45.4 62.5 - 25.3 50.2 62.6 21.5 1 - 105.8
Gold-YOLO-M [39] 49.8 67.0 - 32.3 55.3 66.3 41.3 1 - 197.8
YOLOv8-N [38] 37.2 52.7 40.3 18.9 40.5 52.5 3.2 1 - 20.2
YOLOv8-S [38] 43.9 60.8 47.6 25.3 48.7 59.5 11.2 1 - 66.2
YOLOv8-M [38] 49.8 66.9 54.2 32.6 54.9 65.9 25.9 1 - 182.2
RTMDet-T [55] 41.0 57.4 44.4 20.7 45.3 58.0 4.9 1 - 37.3
RTMDet-S [55] 44.6 61.7 48.3 24.2 49.2 61.8 8.9 1 - 68.1
RTMDet-M [55] 49.3 66.9 53.9 30.5 53.6 66.1 24.7 1 - 180.8
YOLO-MS-XS [8] 42.8 60.0 46.7 23.1 46.8 60.1 5.1 1 - 40.0
YOLO-MS-S [8] 45.4 62.8 49.5 25.9 49.6 62.4 8.7 1 - 69.0
YOLO-MS [8] 49.7 67.2 54.0 32.8 53.8 65.6 23.3 1 - 178.5
YOLOv11-N [40] 39.5 55.3 42.8 19.9 43.3 57.1 2.6 1 - 15.0
YOLOv11-S [40] 47.0 63.9 50.6 29.9 51.6 64.4 9.4 1 - 49.5
YOLOv11-M [40] 51.5 68.4 55.7 33.5 57.0 67.8 20.1 1 - 156.4
YOLOv12-N [62] 40.6 56.7 43.8 20.2 45.2 58.2 2.6 1 - 15.0
YOLOv12-S [62] 48.0 65.0 51.8 30.4 53.2 65.7 9.3 1 - 49.2
YOLOv12-M [62] 52.5 69.6 57.1 35.9 58.2 68.8 20.2 1 - 155.3
SNNs Spiking-YOLO [13] 25.7 - - - - - 10.2 3500 - -
Spike Calib [63] 25.9 - - - - - 10.2 512 - -
EMS-YOLO [15] 30.1 50.1 - - - - 33.9 1$\times$4 - 29.0
M-SpikeFormer [24] - 50.3 - - - - 75.0 4 - 140.8
SUHD [18] - 54.6 - - - - - 4 - -
E-SpikeFormer [21] - 58.8 - - - - 38.7 1$\times$8 - 119.5
SpikeYOLO-N [20] 42.5 59.2 - - - - 13.2 1$\times$4 - 23.1
SpikeYOLO-S [20] 45.5 62.3 49.2 25.5 50.4 61.3 23.1 1$\times$4 12.8 34.6
@@
V-C Comparisons with State-of-The-Art Models
Based on the above ablation studies, we propose SpikeDet-S and SpikeDet-M using MDSNet34 and MDSNet104 as backbones, respectively. Furthermore, to explore the performance upper limit of SNN-based object detectors, we expand the width of SpikeDet-M to propose SpikeDet-L. The comparison results with state-of-the-art methods are shown in Table V. Compared to other SNN-based object detectors, our model achieves significant performance improvements (51.4% AP, 68.6% $\text{AP}_{50}$ vs. previous 48.9% AP, 66.2% $\text{AP}_{50}$, representing +2.5% and +2.4% improvements respectively) while requiring only about $\mathbf{\frac{1}{2}}$ of the power consumption, further strengthening the energy efficiency advantage of SNNs. Furthermore, compared to ANN-based methods, SpikeDet not only further narrows the gap between SNNs and ANNs, but also demonstrates a better trade-off between accuracy and power consumption. For example, SpikeDet-L achieves comparable accuracy to YOLOv12-M while providing 3.6 times the energy efficiency. The inference results of SpikeDet at various scales are presented in Fig. 7.
TABLE VI: Performance Comparison with State-of-the-Art Models on GEN1 Dataset. The $\ddagger$ indicates that the activation function of SpikeDet-S is changed to ReLU, converting it to an ANNs version.
Model AP AP50 Param (M) T×$\times$×D Firing Rate(%) Power (mJ)
RED [64] 40.0 - 24.1 - - $>$24.1
@@ V-C Comparisons with State-of-The-Art Models
ASTMNet [64] 46.7 - $>$100 3 - -
RVT [65] 47.2 - 18.5 10 - 116.7
S5-Vit-B [66] 47.4 - 18.2 10 - -
VC-DenseNet [14] 18.9 - 24.3 5 37.2 5.8
Spiking-YOLO [13] - 45.3 7.9 5 - 0.9
EMS-YOLO [15] 31.0 59.0 14.4 5 17.8 3.4
SFOD [17] 32.1 - 11.9 5 24.4 7.3
EAS-SNN [19] 37.5 69.9 25.3 5 - 28.1
SpikingViT-B [16] 39.4 61.6 21.5 $>$64 18.0 -
SpikeYOLO-S [20] 38.9 66.4 23.1 5$\times$1 - 19.7
SpikeYOLO-S 40.4 67.2 23.1 4$\times$2 - 12.9
SpikeDet-S 46.5 69.2 22.0 5$\times$1 8.8 6.1
SpikeDet-S 47.6 70.1 22.0 4$\times$2 13.8 11.5

### v4

@@ V-A1 Implementation Details
For I-LIF neurons, the membrane time constant $\tau$ is initialized to 0.25. The models are trained on 4 NVIDIA A6000 GPUs, using the SGD optimizer with a learning rate of $0.01$. During post-processing, we use NMS, and the top 100 boxes are selected for evaluation. Additionally, we use Complete IoU (CIOU) [zheng_distanceiou_2020] and Distribution Focal Loss (DFL) [li_generalized_2020] for bounding box regression, and Binary Cross Entropy (BCE) [lin_focal_2017] for classification loss. For multiscale evaluation, following [duan_centernet_2023], we employ both original and horizontally flipped images at scale factors of 0.6, 1.0, 1.2, 1.5, and 1.8.
V-A2 Datasets
@@
The COCO 2017 is a large-scale object detection benchmark with 118,287 training and 5000 validation images, which we use for training and testing, respectively [lin_microsoft_2014]. Objects from 80 categories are annotated with their classes and locations. On this dataset, we train for 400 epochs with a batch size of 64 at a resolution of $\times640$, employing mosaic [lyu_rtmdet_2022] and mixup [zhang_mixup_2018] data augmentation during training.
V-A3 Benchmark Settings
Average Precision. Following [wang_goldyolo_2023, lyu_rtmdet_2022, duan_centernet_2023, zhao_detrs_2024], we report Average Precision (AP) as the primary metric. We also evaluate AP at IoU thresholds of 0.5 and 0.75, as well as AP for small, medium, and large objects.
@@ V-A3 Benchmark Settings
Firing Rate. The firing rate, which measures neuronal activity, is another critical metric for evaluating SNNs. It is calculated as the average ratio of neuron spikes to the total number of neurons across all time steps.
FLOPs and SOPs. We also evaluate the computational cost of ANNs and SNNs. For ANNs, we use FLOPs (FLoating-point OPerations, including additions and multiplications) as the metric. The FLOPs are calculated as:
$\text{FLOPs}=\sum_{n=1}^{N}2\times C^{n-1}\times C^{n}\times(K^{n})^{2}\times H^{n}\times W^{n},$ (16)
where $n$ denotes the layer index, $N$ denotes the total number of layers, $K$ denotes the kernel size, and $C$, $H$, $W$ represent the channel number and spatial dimensions as previously defined. For SNNs, FLOPs is not a suitable metric for computational cost as their computation relies on sparse spike operations without multiplications. Therefore, we employ SOPs (Synaptic OPerations) as the metric, which is calculated as:
$\text{SOPs}=\sum_{t=1}^{T}\sum_{n=1}^{N}{fr}^{t,n-1}\times C^{n-1}\times C^{n}\times(K^{n})^{2}\times H^{n}\times W^{n},$ (17)
where $t$ denotes the time step, $T$ denotes the total number of time steps, and $fr$ denotes the firing rate.
Theoretical Energy Evaluation. Our energy evaluation adopts the standard approach in the SNN field [yao_scaling_2025, luo_integervalued_2025, yao_spikedriven_2024], which employs theoretical computational metrics independent of specific hardware platforms. This theoretical estimation merely serves to facilitate qualitative energy comparisons among different SNN and ANN algorithms. For SNNs, the power consumption is computed as $E_{\text{SNNs}}=E_{\text{AC}}\times\text{ACs}$, where ACs denotes the number of accumulation calculations. Since SOPs consist entirely of accumulation operations, we have $\text{ACs}=\text{SOPs}$. For ANNs, the power consumption is computed as $E_{\text{ANNs}}=E_{\text{MAC}}\times\text{MACs}$, where MACs denotes the number of multiply-accumulate operations, and $\text{FLOPs}=2\times\text{MACs}$. Additionally, following previous works, we adopt $E_{\text{AC}}=0.9\text{pJ}$ and $E_{\text{MAC}}=4.6\text{pJ}$ as the power consumption for a single AC and MAC operation, respectively.
V-B Local Firing Saturation Index
V-B1 Definition
To quantify the effectiveness of our method in addressing the local firing saturation problem, we define a Local Firing Saturation Index (LFSI). For the $n$-th layer with spike feature map $\mathbf{O}\in\mathbb{R}^{T\times C^{n}\times H^{n}\times W^{n}}$, a neuron at $(c,h,w)$ is saturated if it fires at maximum rate:
$\mathbb{I}_{\text{sat}}(\mathbf{O}^{c,h,w})=\begin{cases}1,&\text{if}\sum_{t=1}^{T}\mathbf{O}^{t,c,h,w}=T\times D\\ 0,&\text{otherwise}.\end{cases}$ (18)
For each neuron at location $(c,i,j)$, we compute its local firing saturation density over a $S\times S$ spatial neighborhood $\mathcal{W}_{i,j}$ within channel $c$,
$\rho(c,i,j)=\frac{1}{S^{2}}\sum_{(p,q)\in\mathcal{W}_{i,j}}\mathbb{I}_{\text{sat}}(\mathbf{O}^{c,p,q}).$ (19)
The LFSI for a single layer is the average across all channels and spatial locations,
$\text{LFSI}^{n}=\frac{1}{C^{n}H^{n}W^{n}}\sum_{c=1}^{C^{n}}\sum_{i=1}^{H^{n}}\sum_{j=1}^{W^{n}}\rho(c,i,j).$ (20)
For a network with $N$ spiking layers, the overall LFSI is
$\text{LFSI}=\frac{1}{N}\sum_{n=1}^{N}\text{LFSI}^{n}.$ (21)
A higher LFSI indicates more severe local firing saturation, with $\text{LFSI}\in[0,1]$.
Figure 7: (a) LFSI comparison between SpikeDet and SpikeYOLO across different spatial neighborhood sizes $S$. (b) AP comparison at different IoU thresholds.
Figure 8: Correlation analysis of LFSI across different SNN-based object detectors. (a) Correlation between LFSI and detection performance. (b) Correlation between LFSI and firing rate.
V-B2 Analysis of LFSI Under Varying Spatial Neighborhoods
The parameter $S$ critically affects LFSI evaluation strictness. Small $S$ requires saturated neurons to be spatially contiguous, while large $S$ allows them to be separated within the neighborhood. To investigate this relationship, we analyze LFSI values across different $S$ settings. We conduct experiments on SpikeYOLO [luo_integervalued_2025] and the proposed SpikeDet models to observe the general behavior of this metric. As shown in Fig. 7, the LFSI values of both methods decrease as $S$ increases, confirming the expected relaxation in evaluation criteria. Notably, our method consistently outperforms SpikeYOLO in terms of LFSI across all $S$ values, demonstrating superior capability in mitigating local firing saturation. Moreover, our approach achieves a higher AP across all IoU thresholds, further validating the importance of addressing the local firing saturation problem. Based on these observations, we adopt $S=3$ for subsequent experiments as it provides the most stringent assessment of local firing saturation.
V-B3 Correlation Analysis of LFSI
To further validate the importance of addressing local firing saturation, we analyze the correlation between LFSI and detection performance across different model variants, including SpikeDet, SpikeYOLO, and EMS-YOLO. As illustrated in Fig. 8 (a), LFSI exhibits a strong negative correlation with AP (Pearson r=0.992r=-0.992, $p<0.001$), demonstrating that mitigating local firing saturation is crucial for improving SNN-based object detection. Detailed configurations of SpikeDet variants are provided in Section V-D. Additionally, Fig. 8 (b) reveals a strong positive correlation between LFSI and firing rate (Pearson $r=0.955$, $p<0.001$). Since firing rate directly determines energy consumption as shown in Eq. 17, mitigating local firing saturation effectively reduces power consumption.
V-C Ablation Studies
@@
TABLE I: The ablation study on MDSNet. The specific configuration of the proposed MDSNet is detailed in Appendix B. MS and Plain indicate MDSNet with MDS replaced by membrane-based shortcuts and removed, respectively.
Backbone Depth AP $\text{AP}_{50}$ Param (M) Firing Rate(%) LFSI (%) Power (mJ)
EMS-ResNet [su_deep_2023] 34 43.2 59.5 22.9 14.6 3.2 22.7
SEW-ResNet [fang_deep_2021] 34 43.2 59.7 22.9 12.2 1.6 51.1
MS-ResNet [hu_advancing_2024] 34 43.8 60.5 22.9 12.2 1.7 21.3
S-Backbone [luo_integervalued_2025] 47 44.0 60.6 24.2 12.2 1.7 30.4
MDSNet 34 44.8 61.2 22.0 11.8 1.5 19.0
MDSNet-MS 34 44.2 60.6 21.9 12.0 1.6 19.3
MDSNet-Plain 34 43.7 60.5 21.9 11.4 1.4 18.3
MDSNet 10 41.5 58.2 14.3 12.9 1.9 16.2
MDSNet 18 43.1 59.5 17.5 12.4 1.8 17.5
MDSNet 104 46.5 63.2 48.2 9.1 0.9 23.3
Figure 9: Presynaptic input variance across layers. Experiments are conducted using MDSNet34 and EMS-ResNet34 on COCO. The backbone is divided into 4 stages based on feature map sizes, where I-LIF#1 and I-LIF#2 denote the first and second neurons in the residual path, respectively.
V-C1 The Effectiveness of MDSNet
To demonstrate the feature extraction capabilities of MDSNet, we compare it with the most advanced SNN backbones available. For fair comparison, we construct models using the backbone networks paired with identical neck and detection head architectures. The comparison results are shown in rows 1 to 5 of Table I. These results demonstrate that MDSNet34 significantly outperforms other models in AP and $\text{AP}_{50}$, while achieving lower LFSI, firing rate, and power consumption. Additionally, we conduct comparative experiments by replacing MDS with membrane-based shortcuts and removing MDS. As shown in rows 5 to 7 of Table I, replacing MDS with membrane-based shortcuts increases LFSI, firing rate, and power consumption while reducing AP, demonstrating the effectiveness of MDS in addressing local firing saturation. Removing MDS, despite slightly decreasing LFSI, significantly degrades performance due to the loss of residual learning capability.
To verify the derivation in Section IV-B1, we visualize the presynaptic input variance of EMS-ResNet and MDSNet in Fig. 9. In contrast to EMS-ResNet, our MDSNet maintains stable presynaptic input variance throughout the network. For EMS-ResNet, the presynaptic input to I-LIF#1 (i.e., the output of the previous residual block) exhibits gradually increasing variance, which is consistent with our theoretical analysis in Eq. 12.
V-C2 Ablations on the Depth of MDSNet
In Section IV-C, we theoretically demonstrate that MDSNet overcomes gradient vanishing and exploding, enabling the training of deep networks. In this section, we further validate this through experiments. As shown in rows 5 and 8 to 10 of Table I, the feature extraction capability of MDSNet improves as the network depth increases. Notably, deeper networks exhibit lower LFSI and firing rates, suggesting that deeper architectures can effectively distribute neuronal activity across layers, mitigating local firing saturation issues.
@@ V-C2 Ablations on the Depth of MDSNet
Fusion Method Fusion Directions AP $\text{AP}_{50}$ Param (M) Firing Rate(%) LFSI (%) Power (mJ)
SMFM 1 41.3 58.6 21.1 13.1 1.7 19.5
SMFM 2 43.0 59.9 21.6 12.2 1.6 17.6
SMFM 4 44.8 61.2 22.0 11.8 1.5 19.0
SMFM 6 44.2 60.7 22.1 11.3 1.3 19.6
FPN [lin_feature_2017] 1 38.9 56.4 14.9 14.2 2.2 13.2
PAN [liu_path_2018] 2 39.8 57.7 17.2 12.9 2.0 12.3
BiFPN-D0 [tan_efficientdet_2020] 6 37.5 55.0 11.2 18.2 3.1 8.6
BiFPN-D6 16 42.4 59.9 25.0 11.7 1.1 18.9
S-PAN [luo_integervalued_2025] 2 43.4 60.3 21.0 12.0 1.7 26.4
V-C3 Ablations on Number of Fusion Directions in SMFM
As shown in rows 1 to 4 of Table II, we compare the impact of different fusion direction numbers in SMFM. To ensure fair comparison, we adjust the network width to maintain comparable parameters across models. The results demonstrate that both LFSI and firing rate consistently decrease as the number of fusion directions increases. This validates our analysis in Section IV-D that more fusion directions help mitigate local firing saturation. However, detection performance begins to degrade beyond 4 directions, indicating that excessive structural complexity hinders optimization. Therefore, we adopt 4 fusion directions for SMFM.
We further investigate classical feature fusion modules from both ANN and SNN detectors [lin_feature_2017, liu_path_2018, tan_efficientdet_2020, luo_integervalued_2025], adapting ANN methods for SNNs by replacing activations with I-LIF neurons. As shown in rows 5-9 of Table II, similar trends are observed. More fusion directions consistently lead to better suppression of local firing saturation. Notably, SMFM achieves the best detection performance while maintaining comparable power consumption and lower LFSI and firing rate.
TABLE III: Comparison of different Fusion Blocks. SF-Block- and SF-Block+ refer to removing one MS-Block from SF-Block and adding an additional MS-Block, respectively.
Fusion Block AP $\text{AP}_{50}$ Param (M) Firing Rate(%) LFSI (%) Power (mJ)
$\times7$Conv 41.7 58.5 22.6 11.6 1.9 12.5
MS-Block 43.6 60.7 18.7 13.3 2.0 16.3
SF-Block- 43.8 60.9 20.3 12.0 1.6 17.8
SF-Block+ 44.6 61.2 27.9 11.7 1.5 22.7
SF-Block 44.8 61.2 22.0 11.8 1.5 19.0
V-C4 Ablations on SF-Block
We replace the SF-Block in SMFM with the simplest single convolution to demonstrate its effectiveness. As shown in Table III, SF-Block significantly outperforms it in AP, LFSI, and firing rate, with only a slight increase in power consumption. Meanwhile, we conduct ablation on the two sub-blocks of SF-Block. Rows 2 and 3 in Table III retain only the MS-Block and the first sub-block, respectively, while row 4 extends the MS-Block to two instances. The results demonstrate that the proposed configuration effectively addresses the local firing saturation problem while achieving a better trade-off between performance and power consumption.
V-C5 Ablations on Time Step
Given that SpikeDet provides more stable firing patterns and mitigates local firing saturation, it is expected to achieve comparable performance with fewer time steps. To demonstrate this, we conduct corresponding experiments. Since adjusting T is inefficient [luo_integervalued_2025], we control the overall time steps by adjusting D instead. As shown in Table IV, compared to SpikeYOLO [luo_integervalued_2025], our method achieves good performance even in low time steps. Moreover, performance gains remain modest as the time steps increase, demonstrating stronger robustness to variations in time steps.
@@ V-C5 Ablations on Time Step
TABLE V: Performance Comparison with State-of-the-Art Models on COCO 2017 Validation Dataset. $\dagger$ indicates results from multiscale testing. SNNs include FLOPs in the encoding layer. Therefore, in the computational cost column for SNNs, we report the results in the format of GFLOPs + GSOPs.
@@
ANNs ATSS [zhang_bridging_2020] 39.4 57.6 42.8 23.6 42.9 50.3 32.3 1 - - 164.0 377.2
DETR-DC5-R101 [carion_endtoend_2020] 44.9 64.7 47.7 23.7 49.5 62.3 60.6 1 - - 168.8 388.3
[…]
FCOS [tian_fcos_2019] 44.7 64.1 48.4 27.6 47.5 55.6 89.8 1 - - 459.5 1056.8
TOOD [feng_tood_2021] 46.7 64.6 50.7 28.9 49.6 57.0 51.0 1 - - 272.8 627.4
[…]
GCViTDet [quan2025generalized] 51.8 72.5 56.1 35.1 55.9 67.2 152.0 1 - - 1014.9 2334.2
Gold-YOLO-N [wang_goldyolo_2023] 39.6 55.7 - 19.7 44.1 57.0 5.6 1 - - 12.1 27.6
Gold-YOLO-S [wang_goldyolo_2023] 45.4 62.5 - 25.3 50.2 62.6 21.5 1 - - 46.0 105.8
Gold-YOLO-M [wang_goldyolo_2023] 49.8 67.0 - 32.3 55.3 66.3 41.3 1 - - 87.5 197.8
YOLOv8-N [yolov8_ultralytics] 37.2 52.7 40.3 18.9 40.5 52.5 3.2 1 - - 8.7 20.2
YOLOv8-S [yolov8_ultralytics] 43.9 60.8 47.6 25.3 48.7 59.5 11.2 1 - - 28.6 66.2
YOLOv8-M [yolov8_ultralytics] 49.8 66.9 54.2 32.6 54.9 65.9 25.9 1 - - 78.9 182.2
RTMDet-T [lyu_rtmdet_2022] 41.0 57.4 44.4 20.7 45.3 58.0 4.9 1 - - 8.1 37.3
RTMDet-S [lyu_rtmdet_2022] 44.6 61.7 48.3 24.2 49.2 61.8 8.9 1 - - 14.8 68.1
RTMDet-M [lyu_rtmdet_2022] 49.3 66.9 53.9 30.5 53.6 66.1 24.7 1 - - 39.3 180.8
YOLO-MS-XS [chen_yoloms_2025] 42.8 60.0 46.7 23.1 46.8 60.1 5.1 1 - - 17.4 40.0
YOLO-MS-S [chen_yoloms_2025] 45.4 62.8 49.5 25.9 49.6 62.4 8.7 1 - - 30.0 69.0
YOLO-MS [chen_yoloms_2025] 49.7 67.2 54.0 32.8 53.8 65.6 23.3 1 - - 77.6 178.5
YOLOv11-N [yolo11_ultralytics] 39.5 55.3 42.8 19.9 43.3 57.1 2.6 1 - - 6.5 15.0
YOLOv11-S [yolo11_ultralytics] 47.0 63.9 50.6 29.9 51.6 64.4 9.4 1 - - 21.5 49.5
YOLOv11-M [yolo11_ultralytics] 51.5 68.4 55.7 33.5 57.0 67.8 20.1 1 - - 68.0 156.4
YOLOv12-N [tian_yolov12_2025] 40.6 56.7 43.8 20.2 45.2 58.2 2.6 1 - - 6.5 15.0
YOLOv12-S [tian_yolov12_2025] 48.0 65.0 51.8 30.4 53.2 65.7 9.3 1 - - 21.4 49.2
YOLOv12-M [tian_yolov12_2025] 52.5 69.6 57.1 35.9 58.2 68.8 20.2 1 - - 67.5 155.3
SNNs Spiking-YOLO [kim_spikingyolo_2020] 25.7 - - - - - 10.2 3500 - - - -
EMS-YOLO [su_deep_2023] 30.1 50.1 - - - - 33.9 4 24.9 11.4 1.9 + 27.3 29.0
M-SpikeFormer [yao_spikedriven_2024] - 50.3 - - - - 75.0 1 - - 5.0 + 143.6 140.8
SUHD [qu_spiking_2024] - 54.6 - - - - - 4 - - - -
[…]
V-D Comparisons with State-of-The-Art Models
[…]